OpenStudy (anonymous):

A megabyte is 2^20 bytes. A gigabyte is 2^30 bytes. Using the exponent rule for division, find how many megabytes are in a gigabyte.

ganeshie8 (ganeshie8):

divide

hartnn (hartnn):

divide 2^30 by 2^20 and use \(\huge\frac{a^b }{ a^c }= a^{(b-c)}\)

OpenStudy (kewlgeek555):

I know that since the exponent of ten is positive and it is a six, there is six zeros. 99 should be the first two numbers of the number. So the standard form of 9.9 * 10^6 is 99000000. 99000000 = 9.9 * 10^6 Mrs. Lohr gave me this feedback: Darius, you have an extra 0